OpenStudy (anonymous):

The vertex form of the equation of a parabola is y = 3(x - 4)2 - 22. What is the standard form of the equation?

terenzreignz (terenzreignz):

Evaluate (x - 4)^2 and you'd get...?

OpenStudy (anonymous):

y+22=3(x-4)^2 or (x-4)^2=1/3( y+22) now you can find the vertex of upper parabola by equating to 0 ,(x-4) and (y+22)